OpenStudy (anonymous):

What is the rate of a first-order reaction that has a reactant concentration of 0.1M and a rate constant of 0.1/s?

OpenStudy (jfraser):

what's the general formula for the rate of a first-order reaction?

OpenStudy (anonymous):

r = k[A]^x[B]^y ?

OpenStudy (jfraser):

simplify it to a single reactant \[rate = k[A]^1\] since you know both the rate constant k and the concentration, just plug in to solve for r

OpenStudy (anonymous):

so rate = 0.1M[0.1]^1 ?

OpenStudy (jfraser):

yes

OpenStudy (anonymous):

so 0.01 is the answer, correct?

OpenStudy (jfraser):

yes, but what UNITS does it have? If it's a rate"?

OpenStudy (anonymous):

I honestly have no clue

OpenStudy (jfraser):

the rate constant has units of \(\frac{1}{s}\) and concentration has units of \(\frac{mol}{L}\), so when you multiply them together, you get... \[rate = k*[A]^1 = \frac{1}{s}*\frac{mol}{L} = \frac{mol}{L*s}\]or usually just called "mole per liter per second"